Prof. Finance decides to buy a Nissan Maxima. After paying a down payment and taxes, Prof. Finance can finance the rest of the purchase price with a loan of $28,000 for 60 months at a special finance rate offered by Nissan of 0.9% APR compounded monthly. He finds out that Nissan has a second offer of $2250 cash back (rebate) in place of the special 0.9% finance rate offered. Prof. Finance finds he can get 3.39% APR compounded monthly financing online for 60 months if he takes the $2250 cash back offer. The $2250 cash back would be used as an additional down payment reducing the loan amount.

What is the effective annual rate for each loan option?
Answer #1

Solution: Workings of both the options are as under, Kindly provide necessary feedback with thumbs up or thumbs down.

Decision to Buy Nissan Maxima
Option 1: Borrowing of $28,000.00 for 60 months at 0.9% APR compounded monthly
We would use Present Value Annuity Formula to find the Annuity Payment for this option
We know in our case: Present Value 28000
Time 60
Rate (Compunded) 0.075% 0.9% divided by 12
PVAIF = Annuity*[1-(1+r)^-n]/r
28000= A*48.173
A= =28000/48.173 581.24
Thus monthly repayment in option 1 = $581.24
Effective Rate of Interest = (1+r/no of months in a year)^60 - 1
ERI =(1+0.075%)^60 - 1
4.601%
Option 2: Borrowing of $28,000.00 for 60 months at 3.39% APR compounded monthly with cash back of $2250.00
We would use Present Value Annuity Formula to find the Annuity Payment for this option
We know in our case: Present Value 25750
Time 60
Rate (Compunded) 0.283% 3.39% divided by 12
PVAIF = Annuity*[1-(1+r)^-n]/r
25750= A*55.111
A= =25750/55.111 508.07
Thus monthly repayment in option 2 = $508.07
Effective Rate of Interest = (1+r/no of months in a year)^60 - 1
ERI =(1+0.283%)^60 - 1
18.478%